OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side.

OpenStudy (anonymous):

5n = n + 16 4n = 16 n = 4 Side of the square is 4

OpenStudy (anonymous):

Lord Hades is partially incorrect. n is equal to 4 but the length of the side is 20. (because length = 5n = 5*4)

OpenStudy (anonymous):

Agreed. ^^

OpenStudy (anonymous):

n=4 and one side of the square= 20 in. Don't forget your units!

OpenStudy (anonymous):

It doesn't give the units in the question lol.

OpenStudy (anonymous):

@Lord Hades, yeah it does--the question directly states, "5n inches long," etc.

OpenStudy (anonymous):

Solve the following simultaneous equations for s and n: { s = 5 n, s = n + 16 } s = 20, n = 4